Question 15 0.06 pts Human heights are one of many biological random variables that can be modeled by the normal distribution. A 2002 study in the US found the mean height for 19 year old males is 176.7 cm with a standard deviation of 9.9 cm. A basketball play is usually over 188 cm- what proportion of the 19 year old male population is greater than this height? 1268 O 1306 6268 3732
Answer #1

P(X > 188)

188 -176.7 = P(z >

= P(z > 1.14)

= 0.1268

Option A is correct.